OpenStudy (anonymous):

can -2 be considered a consecutive even integer?

Directrix (directrix):

Consecutive to what other even integer?

OpenStudy (anonymous):

the product of two consecutive even integers is 624. what are the integers?

OpenStudy (anonymous):

I have -24 and positive 26

Directrix (directrix):

-2 and -4 are consecutive even integers -2 and 0 are consecutive even integers.

OpenStudy (anonymous):

so when they ask what is the smaller and larger number, just put -24 smaller and 26 larger?

Directrix (directrix):

@fabo >>so when they ask what is the smaller and larger number, just put -24 smaller and 26 larger? NO

Directrix (directrix):

Let x and (x + 2) be consecutive even integers with a product of 624. If for x, you get -26 AND 24, then you have two sets or pairs of solutions. One set is -26 and (-26 + 2). The other is 24 and (24 + 2). Check them to see if each pair multiplies to 624.

OpenStudy (anonymous):

I get -624

Directrix (directrix):

@fabo What did you get for x? Please post your work. Something is not right.

Directrix (directrix):

(-26)*(-24) = + 624

Directrix (directrix):

The other is 24 and (24 + 2). 24*26 = +624

OpenStudy (anonymous):

(n)(n+2)=624 n^2+2n=624 n^2+2n-624=0 (n+26) (n-24)

OpenStudy (texaschic101):

x = 1st even integer x + 2 = 2nd even integer x(x + 2) = 624 x^2 + 2x = 624 x^2 + 2x - 624 = 0 (x - 24)(x + 26) = 0 x - 24 = 0 x = 24 24 + 2 = 26 24 x 26 = 624 x + 26 = 0 x = -26 -26 + 2 = -24 -26 x -24 = 624

OpenStudy (anonymous):

ok thank you all